# Autonomation, not Automation: Activities and Needs of Fact-checkers as a Basis for Designing Human-Centered AI Systems

### v2

@@ 6.3. Implications for Previously fact-checked claims retrieval
The previously fact-checked claims retrieval has also attracted the researchers’ attention lately, see e.g., (Hossain et al., 2020; Shaar et al., 2021a; Nakov et al., 2022; Mansour et al., 2022; Hardalov et al., 2022). It is typically defined as a ranking problem, where, given an input claim and a set of verified claims (fact-checks), the goal is to retrieve a ranked list of previously fact-checked claims (Shaar et al., 2020). However, the existing datasets and consequently, most of the existing approaches focus on retrieval in a single language - mostly English - although there are also datasets in other languages, namely Arabic (Shaar et al., 2021a) as well as Hindi, Bengali, Malayalam, and Tamil (Kazemi et al., 2021). To the best of our knowledge, only (Kazemi et al., 2021) used multilingual embeddings and provided a preliminary proof-of-concept of multilingual previously fact-checked claims retrieval. Additionally, the existing datasets are relatively small (e.g., the English dataset by (Nakov et al., 2022) contains 1,610 pairs of input claims and fact-checks), they draw input claims from a single social medium (Twitter in the case of (Nakov et al., 2022), WhatsApp in the case of (Kazemi et al., 2021)) and fact-checks from a limited number of fact-checking organizations (mostly Snopes or PolitiFact except of (Kazemi et al., 2021), who use a mixture of sources). The missing cross-lingual links between fact-checks and social media posts in the existing datasets limit their practical applicability since the claims can spread across borders and languages. The research on multilingual and cross-lingual approaches to this task is currently largely missing. This is especially problematic given the results of our study which suggest that searching for other versions of misinformation and searching in existing fact-checks is one of the most important problems among Central and Eastern European fact-checkers and fact-checkers from low-resource languages. The fact-checkers themselves could help foster this type of research by the more widespread use of ClaimReview and especially by providing more appearances of the fact-checked claim (using the ItemReviewed field of the schema). However, this would require better supporting tools and/or training for fact-checkers as discussed in the sections above.
@@
Within this specific AI task, we already did AI research to provide a novel tool for fact-checkers. We collected so far the most extensive and the most linguistically diverse dataset of fact-check to social media post pairs, experimented with and selected the best-performing text embedding models, and built the tool on top of them  (Pikuliak et al., 2023). The whole AI research and tool development process was done in a co-creation manner, in which the identification of fact-checkers’ problems covered by this study played the important first step.
@@ 6.4. Implications for Evidence retrieval
Tools supporting evidence retrieval that would be explainable and trustworthy represent one of the most challenging, but also one of the most urgent needs of fact-checkers. Such tools would need to operate beyond Google searches on websites or extractions from Wikipedia. However, it is exactly these two that are often used when constructing datasets for claim verification; e.g., Wikipedia was used to construct FEVER dataset and its variants (Nørregaard and Derczynski, 2021; Thorne et al., 2018; Ullrich et al., 2022) and Google search was used to retrieve evidence in the case of MultiFC (Augenstein et al., 2019) or X-Fact datasets (Gupta and Srikumar, 2021). Therefore, a more demanding data collection would be crucial for usable and credible verification tools for fact-checkers. One such attempt can be considered the dataset CTKFacts (Ullrich et al., 2022) which extracted evidence from the articles of the Czech Press Agency (CTK), but  databases containing (and combining) the most important sources for fact-checkers, such as the existing fact-checks, official statistics and other official data and the transcriptions of audio-visual content are largely missing. Nevertheless, the introduction of OpenAI’s automatic speech recognition system Whisper has a promising potential for multilingual audio-to-text transcriptions  (Radford et al., 2022).
@@ 6.5. Implications for Claim verification and justification production
The claim verification and justification production, which require a human judgment, was not perceived as problematic by fact-checkers. These professionals enjoyed such challenging tasks requiring higher cognitive load and creativity. Therefore there is no urgency in developing tools that would directly aim to autonomate this stage, and thus support or even replace the fact-checkers. However, the claim verification approaches can be used indirectly as supporting tools to rate or rerank the retrieved evidence and distinguish between sources that support or reject the claim under review (Miranda et al., 2019).
@@ 6.6. Implications for Mapping to additional online content
In the mapping to an additional online content task, the veracity, or knowledge of already fact-checked claims, is automatically disseminated to additional content (already existing or new, constantly emerging). This task is similar to the previously fact-checked retrieval, but, in this case, the input is an already fact-checked claim and the output is a list of claim appearances, i.e., a list of social media posts or news articles, in which the fact-checked claim is present and at the same time, which has a positive stance (they support the claim; see (Srba et al., 2022)). Alternatively, the task can be defined as a combination of previous fact-checked claim retrieval and stance classification. Such an additional step is not usually done in the manual fact-checking process as it is impossible for the fact-checkers to manually find all existing relevant articles or routinely update their list as they continuously appear (Wang et al., 2018). Meta performs this task to some extent – when a new post shares an image that has been previously fact-checked, it propagates the original fact-check to that post. Nevertheless, having an autonomated method that would be able to work also with text in multilingual and cross-lingual settings would further help the fact-checkers increase the impact of their work.
@@ 6.7. Implications for Fact-check summarization and personalization
To the best of our knowledge, this last AI task has not been addressed by existing works so far. The recent emergence and consequent rapid development of generative AI, particularly large language models (LLMs), such as ChatGPT, provides an opportunity to support fact-check dissemination and tackle with fact-checks’ low visibility. Namely, LLMs excel in summarization (Zhang et al., 2023), paraphrasing (Tripto et al., 2023) and translation (Zhu et al., 2023) of the textual content. Therefore, we argue that LLMs can be in the future employed to help fact-checkers process fact-checking articles to shorter and/or more comprehensive forms, such as fact-checking briefs that are suitable to be posted as a social media post. Furthermore, we recognize potential to prepare such LLMs’ prompts that will personalize the fact-checking article for the specific social media platform and target audience.
In this way, fact-checkers can effectively generate a scaffold of the text that can be further manually adjusted (e.g., to clarify potential incomplete information caused by automatic summarization).
@@ 7. Discussion and study limitations
In this study, we first studied the fact-checkers’ activities, needs, and problems. We transformed the findings into practical implications for AI research, AI tasks and AI-based tools. We have presented design implications, which have not yet been used to design and evaluate an AI tool. We defined the stages of the process in a natural setting, the most urgent user needs to autonomize and rejections to do so. Obtaining such results was possible due to the interdisciplinary collaboration (information and computer scientists) as well as due to the involvement of users in the first stages of autonomation. The necessity of such a human-centric AI approach was further demonstrated in the case of the tool for previously fact-checked claim retrieval (Pikuliak et al., 2023). Our research study was the first stage of the HCAI design process, in which end-users are constantly involved in shaping and evaluating the supporting AI tools. Later on, fact-checkers should be actively involved in training AI models by applying a human-in-the-loop approach.
@@
The implications may be valuable not only for AI researchers but also practitioners from technological companies and social media platforms. In general, we recognized that AI has a tremendous potential to help fact-checkers across the whole fact-checking process. Nevertheless, considering the current technical maturity, we suggest using its potential especially in providing support in filtering the monitored content, searching within the existing fact-checks, and evidence retrieval.
@@
Providing multiplatform, multilingual, and multimodal solutions for fact-checkers would be the most useful help for this target group. Our findings extend the implications of (Micallef et al., 2022) that propose personalized multiplatform solutions, including crowdsourcing of user tips; as well as of (Nakov et al., 2021a) who discuss what the technology currently has to offer as well as current major challenges, such as multilingualism and multimodality. The results of our research also pointed to the scenarios that would be ideal, but currently not feasible to autonomate. First, filtering from monitored outlets based on the impact on society would be a task that requires long-term research on how misinformation affects society. Secondly, a tool that would be used globally by fact-checkers would be an ultimate goal for the multinational collaboration and exchange of misinformation narratives, but such a tool would be resource and sustainability heavy as well as context-specific. Thirdly, the ideal content verification would require the availability of (structured) data from governments, which is still not a reality. Fourthly, mapping to additional content to disseminate fact-checks would be impossible without the hands-on cooperation of fact-checkers with social media.
@@
We would also like to acknowledge other limitations of our study. While we aimed to focus specifically on European fact-checkers and complement the existing works, we would like to encourage future researchers to investigate the processes and problems/needs of fact-checkers from the remaining underexplored regions in Asia and Africa. Covering these regions would be beneficial to get a comprehensive global picture of fact-checkers’ routines and problems.
@@
Although the research of the fact-checking activities and problems was directed to the European region, only minor differences were found when comparing these activities and problems with the findings about the fact-checkers from other non-English speaking regions. The implications are also not anyhow specific to individual countries and can be generalized to other contexts with similar attributes (e.g., countries where fact-checkers face the disinformation spread in other languages transmitting from the neighboring countries). This gives the implications a very broad applicability.
@@ 8. Conclusions
In this study, we analyze the activities, processes, and problems of fact-checkers. We focused on European fact-checking space and through this, covered many regions outside of the scope of the existing studies. We compared the details of our research with the findings of existing studies. We have extended the stages of the fact-checking process as it is understood in the current AI research. The hitherto known processes and common problems of this profession were summarized in conceptual models. The fact-checkers’ needs inferred from problems served to inform future AI research on assisting fact-checkers. This is in accordance with the necessity to eliminate the recognized gap between current AI research and fact-checkers’ requirements on and consequent utilization of AI-based tools.
We noted that monitoring the online space to find potential misinformation is one of the most time-consuming and difficult parts of the fact-checking process. As this stage in the fact-checking process is often overlooked in AI research (e.g., (Barrón-Cedeño et al., 2020; Nakov et al., 2021a; Micallef et al., 2022), it is lacking the appropriate support from AI-based information technologies. Besides that, we postulate that without the AI-assisted additional step in the manual fact-checking process mapping to additional online content, any effective mitigation of misinformation is impossible with such a scarce number of fact-checkers.
@@
Fact-checkers do not think that AI can automate their whole work, neither they want it. Fact-checking often requires evaluating complex statements, talking to experts, or determining the truth which are processes inherent to human judgment. Nonetheless, fact-checkers are open and need assistance from AI-based tools, and in this paper, we summarized the possibilities that are offered by the current AI technologies.
Our findings confirm that the identified problems and activities of the target groups can and should serve as potential cases for autonomation to empower, not replace people.
@@ 9. Research ethical considerations
All ideas and content presented in this research are the authors’ own. This study involved human participants and adhered to national and international regulatory laws and ethical standards relevant to such research. Participants were informed that their participation was voluntary and that they could withdraw at any time or request deletion of their data. They were briefed on the research objectives and the use of their data exclusively for research purposes. Additionally, participants were assured that their data would be anonymized, securely stored in the institutional repository, and deleted after publication. While approval from the Institutional Review Board/Ethics Committee was not obtained due to the absence of such a committee at our institution at the time of research design, every effort was made to ensure ethical conduct throughout the study.
@@ II. How big is your institution?
XX. If it was the case, how seriously do you suffer from searching for the image manipulations within the images? Would you appreciate some tech. support (tools) for this work? 5 point Likert scale: 1 = No problem, I like to do it; 5 = I perceive it as a big problem. I really need a help with this
@@
XXVII. Do you wish to make your fact-checks more visible in Google search results or in Google Fact Check Explorer? Example: You would like to use the structure in Schema/ Claim review on your website, but you don’t have any technical capacities or possibilities to do it (even in the Wordpress plugin from Fullfact - Claim review schema) 1) More times a day 2) About once a day 3) About once a week 4) About once a month 5) Less than once a month 6) Never

### v3

@@ 6.3. Implications for Previously fact-checked claims retrieval
The previously fact-checked claims retrieval has also attracted the researchers’ attention lately; see e.g., (Hossain et al., 2020; Shaar et al., 2021a; Nakov et al., 2022; Mansour et al., 2022; Hardalov et al., 2022; Pikuliak et al., 2023; Ramponi et al., 2025). It is typically defined as a ranking problem, where, given an input claim and a set of verified claims (fact-checks), the goal is to retrieve a ranked list of previously fact-checked claims (Shaar et al., 2020). However, the existing datasets and consequently, most of the existing approaches focus on retrieval in a single language – mostly English – although there are also datasets in other languages, such as the one by Kazemi et al. (2021), MultiClaim (Pikuliak et al., 2023), or MMTweets (Singh et al., 2024).
@@
Additionally, the existing datasets often miss cross-lingual links between fact-checks and social media posts. This limits their practical applicability since the claims can spread across borders and languages, which is especially relevant for the European fact-checkers. For example, Pikuliak et al. (2023) acknowledge that there might be as many as seven times more pairs (many of those cross-lingual) in their dataset than what is actually identified. This is especially problematic given the results of our study, which suggest that searching for other versions of misinformation and searching in existing fact-checks is one of the most important problems among Central and Eastern European fact-checkers and fact-checkers from low-resource languages. The fact-checkers themselves could help foster this type of research by the more widespread use of ClaimReview and especially by providing more appearances of the fact-checked claim (using the ItemReviewed field of the schema). However, this would require better supporting tools and/or training for fact-checkers as discussed in the sections above.
The role of LLMs in previously fact-checked claims retrieval is still relatively underexplored (Vykopal et al., 2024), although many solutions to the SemEval 2025 Task 7 task have already used an LLM in their pipeline (Peng et al., 2025). LLMs are currently typically used to rerank the retrieved verified claims (Ramponi et al., 2025), to filter out irrelevant ones (Vykopal et al., 2025), or to decide if the input entails a given verified claim (Choi and Ferrara, 2023). Other options are to employ LLMs as a part of retrieval augmented generation system (RAG) (Lewis et al., 2020), or to use them indirectly, e.g., to normalize claims before retrieval (Sundriyal et al., 2023) or to augment training datasets by synthetic data (Singh et al., 2023).
@@
As part of this AI task, we developed a novel tool for fact-checkers – Fact-check Finder [^26] https://fact-check-finder.kinit.sk/ – using a co-creation process in which identifying fact-checkers’ needs – described in this study – was the essential first step. We have collected so far the most extensive and the most linguistically diverse dataset of fact-checks to social media post pairs, experimented with and selected the best-performing text embedding models, and built the tool on top of them (Pikuliak et al., 2023). Fact-check Finder provides semantic cross-lingual search – an important capability currently limited in Google Fact Check Explorer. For this purpose, Fact-check Finder performs content matching by means of calculating similarity between sentence embeddings created by multilingual language models (Pikuliak et al., 2023).
@@ 6.4. Implications for Evidence retrieval
Tools supporting evidence retrieval that would be explainable and trustworthy represent one of the most challenging but also one of the most urgent needs of European fact-checkers, as evidenced by the results of our study. Such tools would need to operate beyond Google searches on websites or extractions from Wikipedia. However, it is exactly these two that are often used when constructing datasets for claim verification; e.g., Wikipedia was used to construct the FEVER dataset and its variants (Nørregaard and Derczynski, 2021; Thorne et al., 2018; Ullrich et al., 2023), and Google search was used to retrieve evidence in the case of the MultiFC (Augenstein et al., 2019) or X-Fact datasets (Gupta and Srikumar, 2021).
Therefore, a more demanding data collection would be crucial for usable and credible verification tools for fact-checkers. One such attempt can be considered the dataset CTKFacts (Ullrich et al., 2023), which extracted evidence from the articles of the Czech Press Agency (CTK), but databases containing (and combining) the most important sources for fact-checkers, such as official statistics, other official data, and the transcriptions of audio-visual content, are largely missing. Nevertheless, the introduction of OpenAI’s automatic speech recognition system Whisper [^27] https://openai.com/research/whisper has a promising potential for multilingual audio-to-text transcriptions (Radford et al., 2023).
There is also a potential for larger employment of LLMs, especially as a part of retrieval augmented systems (RAGs) (Vykopal et al., 2024), which prompts the LLMs to answer the input based on a set of retrieved examples and/or evidence instead of just relying on the LLMs’ knowledge gained in the pre-training. For example, Momii et al. (2024) employ RAG to retrieve evidence for an input claim by first generating questions that help narrow down the type of evidence needed to verify the input. The evidence can be retrieved from the whole web (as proposed, e.g., in (Ullrich et al., 2024)) as well as from specialized corpora (if available). On the other hand, relying purely on LLMs’ pre-training to provide evidence is not advisable due to possible hallucinations (Huang et al., 2025) and the risks of using a conversational interface to attribute higher credibility to evidence provided in that way (Anderl et al., 2024).
@@ 6.5. Implications for Claim verification and justification production
The claim verification and justification production, which require a human judgment, were not perceived as problematic by the European fact-checkers in our study. These professionals enjoyed such challenging tasks requiring higher cognitive load and creativity. Therefore, there is no urgency in developing tools that would directly aim to autonomate this stage and thus support or even replace the fact-checkers. Despite this fact, automatic claim verification and fact-checking is being actively researched, as also evidenced by an active community around the Fact Extraction and VERification Workshop (FEVER) (Schlichtkrull et al., 2024). We can see the community’s shift towards the use of LLMs and RAGs for verdict production based on the retrieved evidence. While it may not be needed or even preferable due to ethical considerations to use the claim verification approaches directly, they can be used by the fact-checkers indirectly as supporting tools to rate or rerank the retrieved evidence and distinguish between sources that support or reject the claim under review (Miranda et al., 2019). Most recently, Warren et al. (2025) examined the fact-checkers’ requirements regarding the explanations provided by the automated fact-checking systems highlighting the gaps between the current state of the art and the practical needs of fact-checkers.
@@ 6.6. Implications for Mapping to additional online content
In the mapping to an additional online content task, the veracity, or knowledge of already fact-checked claims, is automatically disseminated to additional content (already existing or new, constantly emerging). This task is similar to the previously fact-checked retrieval, but, in this case, the input is an already fact-checked claim, and the output is a list of claim appearances, i.e., a list of social media posts or news articles in which the fact-checked claim is present and at the same time which has a positive stance (they support the claim; see (Srba et al., 2022)). Alternatively, the task can be defined as a combination of previous fact-checked claim retrieval and stance classification. Such an additional step is not usually done in the manual fact-checking process, as it is impossible for the fact-checkers to manually find all existing relevant articles or routinely update their list as they continuously appear (Wang et al., 2018).
At the same time, searching for other versions of misinformation is the second most important problem as identified by the European fact-checkers surveyed in our study. Meta performs (or has performed until recently) this task to some extent – when a new post shares an image that has been previously fact-checked, it propagates the original fact-check to that post. Having an autonomated method that would be able to work also with text in multilingual and cross-lingual settings would further help the fact-checkers increase the impact of their work but also help them monitor how the misinformation spreads and evolves. For this, more research on clustering the misinformation into broader narratives, their identification and evolution is needed (Piskorski et al., 2025).
@@ 6.7. Implications for Fact-check summarization and personalization
The recent emergence and consequent rapid development of generative AI, particularly large language models (LLMs), such as ChatGPT, provides an opportunity to support fact-check dissemination and tackle fact-checks’ low visibility. Namely, LLMs excel in summarization (Zhang et al., 2023), paraphrasing (Tripto et al., 2024) and translation (Zhu et al., 2023) of the textual content. Therefore, we argue that LLMs can be employed to help fact-checkers process fact-checking articles into shorter and/or more comprehensive forms, such as fact-checking briefs that are suitable to be posted as a social media post. A step in this direction presents verdict production, which employs LLMs to generate verdicts (explanations of why a claim is assessed as false or true) by summarizing existing fact-checks (Russo et al., 2023a, b). These verdicts can then be employed on social media, e.g., as responses to social media posts spreading false claims.
Furthermore, we recognize the potential to prepare LLMs’ prompts that will personalize the fact-checking article for the specific social media platform and target audience. In this way, fact-checkers can effectively generate a scaffold of the text that can be further manually adjusted (e.g., to clarify potentially incomplete information caused by automatic summarization). However, it is important to note that for such solutions to be human-centered, they need to respect and follow fact-checking organizations’ guidelines on using LLMs and generative AI to create content.
@@ 7. Discussion and study limitations
In this study, we first studied the fact-checkers’ activities, needs, and problems. We transformed the findings into practical implications for AI research, AI tasks, and AI-based tools. We have presented design implications, which have not yet been used to design and evaluate an AI tool. We defined the stages of the process in a natural setting, the most urgent user needs to autonomate, and rejections to do so. Obtaining such results was possible due to the interdisciplinary collaboration (information and computer scientists) as well as due to the involvement of users in the first stages of autonomation. The practical value of adopting a human-centric AI approach was exemplified by the Fact-check Finder (tool for previously fact-checked claim retrieval (Pikuliak et al., 2023)). Our research study was the first stage of the HCAI design process, in which end-users are constantly involved in shaping and evaluating the supporting AI tools. Later on, fact-checkers should be actively involved in training AI models by applying a human-in-the-loop approach.
@@
The implications in this study may be valuable not only for AI researchers but also for practitioners from technological companies and social media platforms. In general, we recognized that AI has a tremendous potential to help fact-checkers across the whole fact-checking process. Nevertheless, considering the current technical maturity, we suggest using its potential, especially in providing support in filtering the monitored content, searching within the existing fact-checks, and evidence retrieval. An interesting avenue of further research is its use for summarization of fact-checks and their personalization/adaptation for various platforms and/or users.
@@
Providing multiplatform, multilingual, and multimodal solutions for fact-checkers would be the most useful help for this target group. Our findings extend the implications of Micallef et al. (2022), who propose personalized multiplatform solutions, including crowdsourcing of user tips, as well as of Nakov et al. (2021a), who discuss what the technology currently has to offer as well as current major challenges, such as multilingualism and multimodality. The results of our research also pointed to the scenarios that would be ideal but are currently not feasible to autonomate. First, filtering from monitored outlets based on the impact on society would be a task that requires long-term research on how misinformation affects society. Secondly, a tool that would be used globally by fact-checkers would be an ultimate goal for the multinational collaboration and exchange of misinformation narratives, but such a tool would be resource- and sustainability-heavy as well as context-specific. Thirdly, the ideal content verification would require the availability of (structured) data from governments, which is still not a reality. Fourthly, mapping to additional content to disseminate fact-checks would be impossible without the hands-on cooperation of fact-checkers with social media.
@@
We would also like to acknowledge limitations of our study. While we aimed to focus specifically on European fact-checkers and complement the existing works, we would like to encourage future researchers to investigate the processes and problems/needs of fact-checkers from the remaining underexplored regions in Asia and Africa. Covering these regions would be beneficial to get a comprehensive global picture of fact-checkers’ routines and problems.
@@
Although the research of the fact-checking activities and problems was directed to the European region, only minor differences were found when comparing these activities and problems with the findings about the fact-checkers from other non-English-speaking regions. The implications are also not in any way specific to individual countries and can be generalized to other contexts with similar attributes (e.g., countries where fact-checkers face the disinformation spread in other languages transmitting from the neighboring countries). This gives the implications a very broad applicability.
@@ 8. Conclusions
In this study, we analyze the activities, processes, and problems of fact-checkers. We focused on the European fact-checking space, and through this, covered many regions outside of the scope of the existing studies. We compared the details of our research with the findings of existing studies. We have unified the stages of the fact-checking process as it is understood in the current social-science and AI research. The hitherto known processes and common problems of this profession were summarized in conceptual models. The fact-checkers’ needs inferred from problems served to inform future AI research on assisting fact-checkers. This is in accordance with the necessity to eliminate the recognized gap between current AI research and fact-checkers’ requirements on and consequent utilization of AI-based tools.
We noted that monitoring the online space to find potential misinformation is one of the most time-consuming and difficult parts of the fact-checking process. As this stage in the fact-checking process is often overlooked in AI research (e.g., (Barrón-Cedeño et al., 2020; Nakov et al., 2021a; Micallef et al., 2022), it is lacking the appropriate support from AI-based information technologies. Besides that, we postulate that without the AI-assisted additional step in the manual fact-checking process – mapping to additional online content – any effective mitigation of misinformation is impossible with such a scarce number of fact-checkers.
@@
Fact-checkers do not think that AI can automate their whole work, nor do they want it. Fact-checking often requires evaluating complex statements, talking to experts, or determining the truth, which are processes inherent to human judgment. Nonetheless, fact-checkers are open and need assistance from AI-based tools, and in this paper, we summarized the possibilities that are offered by the current AI technologies.
Our findings confirm that the identified problems and activities of the target groups can and should serve as potential cases for autonomation to empower, not replace, people.
@@ 9. Research ethical considerations
All ideas and content presented in this research are the authors’ own. This research involved human participants. Research planning, conduct, and reporting were consistent with local regulatory laws and regulations and aligned with ethical principles, such as the ACM Code of Ethics and Professional Conduct and international and national standards for such research. Participants were informed that their participation was voluntary and that they could withdraw at any time or request deletion of their data. They were briefed on the research objectives and the use of their data exclusively for research purposes. Additionally, participants were assured that their data would be anonymized, securely stored in the institutional repository, and deleted after publication. While approval from the Institutional Review Board/Ethics Committee was not obtained due to the absence of such a committee at our institution at the time of research design, every effort was made to ensure ethical conduct throughout the study.
Acknowledgements.
This work was partially supported by CEDMO, a project funded by the European Union under the Contract No. 2020-EU-IA-0267; CEDMO 2.0, a project funded by the European Union under the Contract No. 101158609; AI-CODE, a project funded by the European Union under the Horizon Europe, GA No. 101135437; and by the EU NextGenerationEU through the Recovery and Resilience Plan for Slovakia under the project No. 09I01-03-V04-00006.
@@ II. How big is your institution?
$X$. If it was the case, how seriously do you suffer from searching for the image manipulations within the images? Would you appreciate some tech. support (tools) for this work? 5 point Likert scale: 1 = No problem, I like to do it; 5 = I perceive it as a big problem. I really need a help with this
@@
XXVII. Do you wish to make your fact-checks more visible in Google search results or in Google Fact Check Explorer?      Example: You would like to use the structure in ClaimReview Schema on your website, but you don’t have any technical capacities or possibilities to do it (even in the Wordpress plugin from Fullfact – ClaimReview Schema) 1) More times a day 2) About once a day 3) About once a week 4) About once a month 5) Less than once a month 6) Never